# Signal-to-noise Ratio Analytic Formulae of the Inspiral Binary Black Holes in TianQin

### v2

@@ Abstract
Binary black holes are one of the important sources for the TianQin gravitational wave project. Our research has revealed that, for TianQin, the signal-to-noise ratio of inspiral binary black holes can be computed analytically. This finding is expected to greatly simplify the estimation of detection capabilities for binary black holes. In this paper, we demonstrated the signal-to-noise ratio relationships from stellar-mass black holes to massive black holes. With the all-sky average condition, the signal-to-noise ratio for most binary black hole signals can be determined with a relative error of $\lesssim 10\%$, with notable deviations only for chirp masses near $1000~{}M_{\odot}$. In contrast, the signal-to-noise ratio without the average includes an additional term, which we refer to as the response factor. Although this term is not easily calculated analytically, we provide a straightforward estimation method with an error margin of $1\sigma$ within 2%.
preprint: APS/123-QED
I Introduction
In recent years, the domain of gravitational wave (GW) detection has undergone remarkable advancements, marking a significant milestone in the scientific community. One of the key breakthroughs came in 2015, when Laser Interferometer GW Observatory (LIGO) [1] successfully detected the first GW signal from a binary black hole (BBH) merger [2, 3]. So far, nearly a hundred of GW detections have been officially announced by ground-based detectors [4, 5, 6]. Furthermore, evidence for the stochastic GW background has recently been reported by pulsar timing array (PTA) [7, 8, 9, 10]. In parallel, research on space-based GW detectors is also progressing [11]. For example, the TianQin project [12], a space-borne GW detection mission, aims to detect low-frequency GWs and further advance our understanding of the universe[13].
@@ I Introduction
For SBBHs, they merge at high frequencies, space-based GW detectors can observe their early inspiral phase within the lower frequency range. The detection of SBBHs in the millihertz band with a long observation time enables more precise measurements of parameters such as mass and sky location, while also preserving the evolution of key properties such as eccentricity and spin [22, 23]. These capabilities not only help constrain the formation mechanisms of such systems [24, 25], but also facilitate measuring the Hubble parameter or constraining certain parameters in modified gravity theories with great precision [26, 27, 28, 29]. We have found that, for TianQin, the SNR of GWs emitted by inspiral SBBHs follows a special relationship:
@@
These relationships will significantly facilitate the estimation of the SNR for the inspiral phase of BBHs, as well as the forecast of detection capabilities before coalescence. In this work, we will conduct a more detailed derivation and discussion of these relationships. Section II will outline the formula for calculating the SNR. Sections III explore the SNR accumulation for MBHB systems in the TianQin detector, under All-Sky Average (ASA) and non-average conditions respectively. Then, Sections IV delves into the feasibility of formulating an analytical SNR formula for the inspiral SBBHs. We conclude with a summary in Section V.
II SNR calculation
@@
where $S_{n}\left(f\right)$ is the one-sided noise power spectral density (PSD) and $\Re$ is the real component. The instrument settings used for the TianQin/LISA PSD are detailed in Appendix A.
@@
Here, $G$ and $c$ denote the gravitational constant and the speed of light, $t_{c}$ and $\mathcal{M}$ correspond to the coalescence time and the detector-frame chirp mass. Setting the time origin at the onset of the observation period, the upper and lower limits of the frequency of the SNR integration are $f_{\rm max}=f\left(T_{\rm obs}\right)$ and $f_{\rm min}=f(0)$, respectively. Moreover, we need to cutoff the data to reduce errors if the signal frequency exceeds the cutoff frequency of TianQin. A detailed description can be found in the Appendix A.
@@
III Massive Black Holes
III.1 All-Sky Average
By using low-frequency approximation, we can provide the analytic formula of the SNR for the TianQin detector’s observation of inspiral MBHBs:
$\rho(T_{\rm obs})$ $=\sqrt{\frac{15}{2048}}\frac{L}{D_{L}}\frac{c}{\sqrt{N_{a}t_{c}}}\times\sqrt{\frac{t_{obs}}{t_{c}-t_{obs}}},$
$\approx 185.2\times\frac{1~{}{\rm Gpc}}{D_{L}}\sqrt{\frac{1~{}{\rm week}}{t_{c}}}\times\sqrt{\frac{T_{\rm obs}}{t_{c}-T_{\rm obs}}},$ (7)
A detailed description of the calculation process can be found in the Appendix B.
To verify the precision of the estimation formula, we examined the error levels across different chirp masses and observation times. In our test, all the BBH systems are merged at the third month mark $\left(t_{c}=3~{}{\rm months}\right)$, and the true signal is generated by IMRPhenomD waveform [34, 35]. As depicted in Figure 1(a), for signals with chirp mass range from $10^{3}$ to $10^{4}M_{\odot}$, the relative error is minimal. Signals outside this region will be overestimated by our formula. Considering the low SNR of the inspiral BBHs, the deviation caused by this relative error will be very low, ensuring the reliability of the formula.
@@ III.1 All-Sky Average
$\rho_{\rm corr}^{2}=\rho^{2}+C,$ (8)
where
$C_{TQ}=$ $-\frac{3\pi c^{1/8}}{320000\times 5^{3/8}}\left(G\mathcal{M}\right)^{5/8}\left(\frac{L}{D_{L}}\right)^{2}\frac{1}{N_{a}}$
×$\times\left[\left(t_{c}-t_{{\rm obs}}\right)^{-5/8}-t_{c}^{-5/8}\right],$
$\approx$ $-1692\left(\frac{\mathcal{M}}{10^{4}~{}M_{\odot}}\right)^{5/8}\left(\frac{1~{}{\rm Gpc}}{D_{L}}\right)^{2}$
×$\times\left[\left(\frac{t_{c}-T_{\rm obs}}{1~{}{\rm week}}\right)^{-5/8}-\left(\frac{t_{c}}{1~{}{\rm week}}\right)^{-5/8}\right].$ (9)
Figure 1(b) illustrates the improvement brought by this correction. Compared to Figure 1(a), the corrected SNR formula provides an accurate estimation for sources within the mass range of $10^{3}$ to $10^{5}M_{\odot}$. Signals with masses falling below this range are prone to be overestimated by our formula, whereas signals with masses exceeding this range are susceptible to underestimation. Systems with $\mathcal{M}\leq 10^{3}M_{\odot}$ retain relatively poor estimation because the low-frequency approximation we used here is no longer suitable for these systems.
@@
Figure 1: Estimated error of the SNR for signals that merge at the third month, under different observation times and chirp mass conditions. The left and right panels respectively illustrate the estimated errors of Equation III.1 and (8). The blank area in the lower right corner indicates that the signal has not yet entered the sensitive frequency band of TianQin.
We’ve also noted that this formula might not be well-suited for other space-based GW detectors, such as Laser Interferometer Space Antenna (LISA) [36]. LISA, which will be placed in heliocentric orbit, consists of three spacecraft forming an equilateral triangle with each arm spanning 2.5 million kilometers. The ASA sensitivity curve of LISA aligns with that of TianQin, differing only in arm length and two noise terms, as detailed in Appendix A. The increase in arm length enhances sensitivity in the low-frequency band, prompting us to select a lower cutoff frequency for LISA at $10^{-5}$ Hz rather than $10^{-4}$ for TianQin. Different dependencies of $S_{a}$ on frequency only alter the higher-order terms of the SNR formula. For LISA, the lowest-order correction is:
$C_{LISA}=$ $-\frac{3\pi^{2}}{2\times 10^{6}\times(5c)^{7/4}}\left(G\mathcal{M}\right)^{5/4}\left(\frac{L}{D_{L}}\right)^{2}\frac{1}{N_{a}}$
×$\times\left[\left(t_{c}-T_{\rm obs}\right)^{-1/4}-t_{c}^{-1/4}\right],$
$\approx$ $-48298\left(\frac{\mathcal{M}}{10^{4}~{}M_{\odot}}\right)^{5/4}\left(\frac{1~{}{\rm Gpc}}{D_{L}}\right)^{2}$
×$\times\left[\left(\frac{t_{c}-T_{\rm obs}}{1~{}{\rm week}}\right)^{-1/4}-\left(\frac{t_{c}}{1~{}{\rm week}}\right)^{-1/4}\right]$ (10)
However, as shown in Figure 2, the estimated results of the SNR formula always have significant errors, regardless of whether the lowest-order correction is taken into account. The main reason is that the higher-order corrections in the LISA SNR formula are not negligible quantities. When not considering the lowest-order correction, as shown in Figure 2(a), the estimated error remains above 10%. Even with the lowest-order correction, only sources with masses around $3\times 10^{4}M_{\odot}$ can be estimated accurately. Furthermore, we attempted to vary the cutoff frequencies for LISA, with the low-frequency cutoff ranging from $10^{-5}$ Hz to $10^{-4}$ Hz, and the high-frequency cutoff ranging from 0.1 Hz to 1 Hz. Our findings indicate that there is no substantial difference in the magnitude of estimation errors.
@@ III.2 Non-average
The expression $P_{ab}=n_{l}\cdot P^{22}\cdot n_{l}$ defines the inner product of the 2-2 mode polarization tensor $P^{22}$ and the link unit vectors $n_{l}$ between two satellites $a$ and $b$. A detailed description of the calculation process can be found in the Appendix C.
Figure 3 illustrates the variation of the response factor across different sky positions and polar angles. Given the response factor’s dependence on all four positional angles, we integrate the remaining two angles to ascertain their relationship with a pair of selected angles during the computation process. As shown in Figure 3(a), the TianQin constellation exhibits pronounced sensitivity towards the double white dwarf system RX J0806.3 + 1527 (hereafter J0806) and the reciprocal direction, attributable to its orbital plane’s always facing the J0806. The orbital plane of TianQin and the coordinates of J0806 in the ecliptic reference frame are marked using red dashed line and red asterisk, respectively. In Figure 3(b), it is evident that the inclination angle $\iota$ has a more significant impact on the SNR compared to the polarization angle $\psi$. The SNR is maximized when the inclination angle approaches 00 or $\pi$, indicating that the detector is face-on the source. Conversely, the SNR is minimal when the inclination angle is near $\pi/2$, and it is under this condition that the polarization angle exerts a noticeable influence on the SNR. In contrast, at the poles of Figure 3(b), the contour lines are almost coincident with the lines of latitude, indicating that the influence of the polarization angle $\psi$ can be disregarded in this area.
@@
$P_{res}=P(\lambda,\beta)\times P(\psi,\iota).$ (13)
@@
IV Stellar-mass Black Holes
Besides MBHBs, we are further exploring the feasibility of deriving an analytical SNR formula for inspiral SBBHs. Due to the relatively high frequency at which SBBHs evolves within TianQin’s sensitive band, the low-frequency approximation is no longer applicable. Conversely, the PSD formula of TianQin in ASA case can be simplified as follows:
@@ IV Stellar-mass Black Holes
Using the same methodology of MBHB, we can obtain the SNR:
$\rho(T_{\rm obs})=$ $\frac{1.15829}{c^{11/4}\sqrt{N_{p}}}\left(G\mathcal{M}\right)^{5/4}\frac{L}{D_{L}}\times\left(\sqrt{t_{c}}-\sqrt{t_{c}-T_{\rm obs}}\right)^{1/2}$
$\approx$ $8.4\left(\frac{\mathcal{M}}{100~{}M_{\odot}}\right)^{5/4}\frac{100~{}{\rm Mpc}}{D_{L}}$
×$\times\left(\sqrt{\frac{t_{c}}{\rm 1~{}month}}-\sqrt{\frac{t_{c}-T_{\rm obs}}{\rm 1~{}month}}\right)^{1/2},$ (15)
@@
$C=$ $-\frac{0.388101}{c^{13/3}N_{p}}\left(G\mathcal{M}\right)^{5/3}\frac{L^{10/3}}{D_{L}^{2}}\times\left[\mathcal{F}(t_{c})-\mathcal{F}(t_{c}-T_{\rm obs})\right],$
$\approx$ $162\left(\frac{\mathcal{M}}{100~{}M_{\odot}}\right)^{5/3}\left(\frac{100~{}{\rm Mpc}}{D_{L}}\right)^{2}\times\left[\mathcal{F}(t_{c})-\mathcal{F}(t_{c}-T_{\rm obs})\right],$ (16)
@@
$\mathcal{F}(t)=$ ${\rm arctan}\left[0.57735-\frac{0.577948~{}c^{7/12}L^{2/3}}{\left(G\mathcal{M}\right)^{5/12}t^{1/4}}\right],$
$\approx$ ${\rm arctan}\left[0.57735-\frac{0.23866}{\left(\frac{\mathcal{M}}{100~{}M_{\odot}}\right)^{5/12}\left(\frac{t}{\rm 1~{}month}\right)^{1/4}}\right]$ (17)
In our test, as shown in Figure 5, we found that in the case of ASA, our formula provided good estimation accuracy for SBBHs. Without incorporating the lowest-order correction, we could make good estimates of the SNR for signals ranging from $5\sim 500~{}M_{\odot}$. The SNR was overestimated for signals outside this range. When the lowest-order correction was applied, the SNR of most signals with masses below $500~{}M_{\odot}$ could be accurately calculated.
@@
Despite our formula providing good estimations in the case of ASA, the SBBH signal with response is challenging to compute analytically. As illustrated in Figure 6, after considering the TDI response, the low-frequency component of the PSD remains smooth, allowing the SNR of the MBHB signal with response to be analytically approximated using low-frequency approximation. While in the high-frequency regime where SBBH resides, the PSD becomes highly complex and lacks a satisfactory formula for approximation. However, it is important to note that while we consider a three-month observation for all systems, SBBHs can keep inspiraling within the sensitive bands of space detectors for years. Therefore, our estimation of the ASA SNR will also be applicable for such long-duration inspiral-only observations.
@@
V Conclusion
In this work, we present a simplified analytic formula for the SNR of BBH systems in the context of TianQin. The analytical formula for low-frequency scenarios is applicable to signals from MBHBs with chirp masses above approximately $10^{3}~{}M_{\odot}$, while the analytical formula for high-frequency scenarios is suitable for signals from SBBHs with chirp masses below approximately $500~{}M_{\odot}$
We evaluated the significance of errors resulting from using SNR analytical formulae. For TianQin, our formula of MBHB can accurately estimate the inspiral signals within the mass range of $10^{3}$ to $10^{4}~{}M_{\odot}$. Incorporating a lowest-order correction can extend the range of accurate estimates between $10^{3}$ and $10^{5}~{}M_{\odot}$. However, for LISA, achieving a good estimation accuracy is challenging, regardless of whether the lowest-order correction is considered.
Then, we derived an accurate analytical formula for the SNR of SBBHs in the ASA case. The formula can accurately estimate the inspiral signals within the mass range of $5$ to $500~{}M_{\odot}$. Upon the application of the lowest-order correction, the SNR of the majority of signals with chirp mass below $500~{}M_{\odot}$ could be accurately calculated. Due to the complexity of the PSD at high frequencies, it remains challenging to provide a precise analytical formula for the SNR of SBBH signals with response. We will delve into this scenario in the future.

### v3

@@ Abstract
Binary black holes are one of the important sources for the TianQin gravitational wave project. Our research has revealed that, for TianQin, the signal-to-noise ratio of inspiral binary black holes can be computed analytically. This finding will help to quickly estimate the detectability of gravitational waves from different types of binary black holes. In addition, it provides us with the accumulation pattern of signal-to-noise ratios for binary black hole signals, which can be used for the identification of future signals and the construction of detection statistics. In this paper, we demonstrated the signal-to-noise ratio analytic formulae from stellar-mass black holes to massive black holes. With the all-sky average condition, the signal-to-noise ratio for most binary black hole signals can be determined with a relative error of $\lesssim 10\%$, with notable deviations only for chirp masses near $1000~{}M_{\odot}$. Moreover, the signal-to-noise ratio without the average includes only an additional coefficient, which we refer to as the response factor. Although this term is not easily calculated analytically, we provide a straightforward calculation method with an error margin of $1\sigma$ within 2%.
1 Introduction
In recent years, the domain of gravitational wave (GW) detection has undergone remarkable advancements, marking a significant milestone in the scientific community. One of the key breakthroughs came in 2015, when Laser Interferometer GW Observatory (LIGO) [1] successfully detected the first GW signal from a binary black hole (BBH) merger [2, 3]. So far, hundreds of GW detections have been made by ground-based detectors [4, 5, 6, 7]. Furthermore, evidence for the stochastic GW background has recently been reported by pulsar timing array (PTA) [8, 9, 10, 11]. In parallel, research on space-based GW detectors is also progressing [12]. For example, the TianQin project [13], a space-borne GW detection mission, aims to detect low-frequency GWs and further advance our understanding of the universe [14].
@@ 1 Introduction
For SBBHs, they merge at high frequencies. Space-based GW detectors can observe their early inspiral phase within the lower frequency range, with the SNR reaching over a dozen [23, 24]. The detection of SBBHs in the millihertz band with a long observation time enables more precise measurements of parameters such as mass and sky location, while also preserving the evolution of key properties such as eccentricity and spin [23, 24]. These capabilities not only help constrain the formation mechanisms of such systems [25, 26], but also facilitate measuring the Hubble parameter or constraining certain parameters in modified gravity theories with great precision [27, 28, 29, 30]. We have found that, for TianQin, the SNR of GWs emitted by inspiral SBBHs follows a special relationship:
@@
These relationships will significantly facilitate the estimation of the SNR for the inspiral phase of BBHs, simplify the assessment of the detectability of different BBH models, and be beneficial for the identification of future signals and the construction of detection statistics. In this paper, we first present a general SNR calculation method for the inspiral phase. Then, we derive the SNR relationships for MBHB and SBBH using Low-Frequency Approximation (LFA) and Short-Wave Approximation (SWA) approximations respectively. For MBHB, we study the inspiral SNR relationships under both All-Sky Average (ASA) (Equation 7) and non-average (Equation 13) conditions. For SBBH, we only focus on the inspiral SNR relationship under ASA (Equation 17).
The organization of this paper is as follows. Section 2 will outline the formula for calculating the SNR. Section 3 explores the SNR accumulation for MBHB systems in the TianQin detector, under ASA and non-average conditions respectively. Then, Section 4 delves into the feasibility of formulating an analytical SNR formula for the inspiral SBBHs. We conclude with a summary in Section 5.
2 SNR calculation
@@
where $S_{n}\left(f\right)$ is the one-sided noise power spectral density (PSD) and $\Re$ is the real component. The instrument settings used for the TianQin/LISA PSD are detailed in A.
@@
Here, $G$ and $c$ denote the gravitational constant and the speed of light, $t_{c}$ and $\mathcal{M}$ correspond to the coalescence time and the detector-frame chirp mass. Setting the time origin at the onset of the observation period, the upper and lower limits of the frequency of the SNR integration are $f_{\rm max}=f\left(T_{\rm obs}\right)$ and $f_{\rm min}=f(0)$, respectively. Moreover, we need to cutoff the data to reduce errors if the signal frequency exceeds the cutoff frequency of TianQin. A detailed description can be found in the A.
@@
3 Massive Black Holes
3.1 All-Sky Average
By using the LFA, we can provide the analytic formula of the SNR for the TianQin detector’s observation of inspiral MBHBs:
$\rho(T_{\rm obs})$ $=\sqrt{\frac{15}{2048}}\frac{L}{D_{L}}\frac{c}{\sqrt{N_{a}t_{c}}}\times\sqrt{\frac{t_{obs}}{t_{c}-t_{obs}}},$ (7)
$\approx 185.2\times\frac{1~{}{\rm Gpc}}{D_{L}}\sqrt{\frac{1~{}{\rm week}}{t_{c}}}\times\sqrt{\frac{T_{\rm obs}}{t_{c}-T_{\rm obs}}},$
A detailed description of the calculation process can be found in the B. It is worth noting that the formula presented here does not include the chirp mass parameter. This indicates that due to TianQin’s frequency power-law sensitivity curve, in specific frequency bands, variations in chirp mass do not significantly affect the SNR of MBHB GW signals. This observation aligns with findings in previous work [22].
If there is a gap in the data, the SNR can be expressed as:
$\rho=\sqrt{\rho^{2}\left(T_{\rm obs}\right)-\rho^{2}\left(t_{e}\right)+\rho^{2}\left(t_{s}\right)},$ (8)
where, $t_{s}$ and $t_{e}$ indicate the start and end times of the gap. More gaps can also be calculated simply using this similar method.
To verify the precision of the estimation formula, we examined the error levels across different chirp masses and observation times. In our test, the true signal is generated by the IMRPhenomD waveform [35, 36], and all the BBH systems are merged at the third month mark $\left(t_{c}=3~{}{\rm months}\right)$, as it illustrates the possible maximum observation time in the future. We define the relative error as
$\Delta=\frac{\rho-\rho_{\rm true}}{\rho_{\rm true}}.$ (9)
As depicted in Figure 1(a), for signals with chirp mass range from $10^{3}$ to $10^{4}M_{\odot}$, the relative error is minimal. Signals outside this region will be overestimated by our formula. Considering the low SNR of the inspiral BBHs, the deviation caused by this relative error will be very low, ensuring the reliability of the formula.
@@ 3.1 All-Sky Average
$\rho_{\rm corr}^{2}=\rho^{2}+C.$ (10)
The details of the calculation method for the correction term are provided in the C. For TianQin,
$C_{TQ}$ $=-\frac{3\pi c^{1/8}}{320000\times 5^{3/8}}\left(G\mathcal{M}\right)^{5/8}\left(\frac{L}{D_{L}}\right)^{2}\frac{1}{N_{a}}\times\left[\left(t_{c}-t_{{\rm obs}}\right)^{-5/8}-t_{c}^{-5/8}\right],$ (11)
$\approx-1692\left(\frac{\mathcal{M}}{10^{4}~{}M_{\odot}}\right)^{5/8}\left(\frac{1~{}{\rm Gpc}}{D_{L}}\right)^{2}\times\left[\left(\frac{t_{c}-T_{\rm obs}}{1~{}{\rm week}}\right)^{-5/8}-\left(\frac{t_{c}}{1~{}{\rm week}}\right)^{-5/8}\right].$
Figure 1(b) illustrates the improvement brought by this correction. Compared to Figure 1(a), the corrected SNR formula provides an accurate estimation for sources within the mass range of $2\times 10^{3}\sim 2\times 10^{5}M_{\odot}$. Signals with masses falling below this range are prone to be overestimated by our formula, whereas signals with masses exceeding this range are susceptible to underestimation. Systems with $\mathcal{M}\leq 10^{3}M_{\odot}$ retain relatively poor estimation because the LFA we used here is no longer suitable for these systems.
@@
Figure 1: Estimated error of the SNR for signals that merge at the third month, under different observation times and chirp mass conditions. The left and right panels respectively illustrate the estimated errors of Equation 7 and (10). The blank area in the lower right corner indicates that the signal has not yet entered the sensitive frequency band of TianQin.
We’ve also noted that this formula might not be well-suited for other space-based GW detectors, such as Laser Interferometer Space Antenna (LISA) [37]. LISA, which will be placed in heliocentric orbit, consists of three spacecraft forming an equilateral triangle with each arm spanning 2.5 million kilometers. The ASA sensitivity curve of LISA aligns with that of TianQin, differing only in arm length and two noise terms, as detailed in A. The increase in arm length enhances sensitivity in the low-frequency band, prompting us to select a lower cutoff frequency for LISA at $10^{-5}$ Hz rather than $10^{-4}$ for TianQin. Different dependencies of $S_{a}$ on frequency only alter the higher-order terms of the SNR formula. For LISA, the lowest-order correction is:
$C_{LISA}$ $=-\frac{3\pi^{2}}{2\times 10^{6}\times(5c)^{7/4}}\left(G\mathcal{M}\right)^{5/4}\left(\frac{L}{D_{L}}\right)^{2}\frac{1}{N_{a}}\times\left[\left(t_{c}-T_{\rm obs}\right)^{-1/4}-t_{c}^{-1/4}\right],$ (12)
$\approx-48298\left(\frac{\mathcal{M}}{10^{4}~{}M_{\odot}}\right)^{5/4}\left(\frac{1~{}{\rm Gpc}}{D_{L}}\right)^{2}\times\left[\left(\frac{t_{c}-T_{\rm obs}}{1~{}{\rm week}}\right)^{-1/4}-\left(\frac{t_{c}}{1~{}{\rm week}}\right)^{-1/4}\right]$
However, as shown in Figure 2, the estimated results of the SNR formula always have significant errors, regardless of whether the lowest-order correction is taken into account. Due to experiments like LISA Pathfinder, LISA has gained a more profound understanding of noise characteristics. Consequently, it has developed a more complex formula for the acceleration noise power spectrum. Applying our simplified approaches to LISA may introduce larger errors. However, such a simplification will not result in significant errors for TianQin with a relatively simpler noise design curve. When not considering the lowest-order correction, as shown in Figure 2(a), the estimated error remains above 10%. Even with the lowest-order correction, only sources with masses around $3\times 10^{4}M_{\odot}$ can be estimated accurately. Other LISA-like missions (e.g., Taiji) have similar orbits and noise power spectral density, so their SNR formulas’ structures and applicability are basically the same as LISA’s. Furthermore, we attempted to vary the cutoff frequencies for LISA, with the low-frequency cutoff ranging from $10^{-5}$ Hz to $10^{-4}$ Hz, and the high-frequency cutoff ranging from 0.1 Hz to 1 Hz. Our findings indicate that there is no substantial difference in the magnitude of estimation errors.
@@ 3.2 Non-average
The expression $P_{ab}=n_{l}\cdot P^{22}\cdot n_{l}$ defines the inner product of the 2-2 mode polarization tensor $P^{22}$ and the link unit vectors $n_{l}$ between two satellites $a$ and $b$. Note that this concept is distinct from the antenna pattern and focuses on the ratio of the response at a specific sky position and polar angle to the ASA. A detailed description of the calculation process can be found in the D. And we offer a simple and quick calculation program in https://github.com/TianQinSYSU/responseFactor.
Figure 3 illustrates the variation of the response factor across different sky positions and polar angles. Given the response factor’s dependence on all four positional angles, we integrate the remaining two angles to ascertain their relationship with a pair of selected angles during the computation process. As shown in Figure 3(a), the TianQin constellation exhibits pronounced sensitivity towards the double white dwarf system RX J0806.3 + 1527 (hereafter J0806) and the reciprocal direction, attributable to its orbital plane’s always facing the J0806. The orbital plane of TianQin and the coordinates of J0806 in the ecliptic reference frame are marked using a red dashed line and a red asterisk, respectively. In Figure 3(b), it is evident that the inclination angle $\iota$ has a more significant impact on the SNR compared to the polarization angle $\psi$. The SNR is maximized when the inclination angle approaches 00 or $\pi$, indicating that the detector is face-on to the source. Conversely, the SNR is minimal when the inclination angle is near $\pi/2$, and it is under this condition that the polarization angle exerts a noticeable influence on the SNR. In contrast, at the poles of Figure 3(b), the contour lines are almost coincident with the lines of latitude, indicating that the influence of the polarization angle $\psi$ can be disregarded in this area.
@@
$P_{\rm res}=P(\lambda,\beta)\times P(\psi,\iota).$ (15)
@@
4 Stellar-mass Black Holes
Besides MBHBs, we are further exploring the feasibility of deriving an analytical SNR formula for inspiral SBBHs. Due to the relatively high frequency at which SBBHs evolves within TianQin’s sensitive band, the LFA is no longer applicable. Conversely, the PSD formula of TianQin in ASA case can be simplified by SWA as follows:
@@ 4 Stellar-mass Black Holes
Using the same methodology of MBHB, we can expand the SNR integral and retain only the zero-order term.:
$\rho(T_{\rm obs})$ $=$ $\frac{1.15829}{c^{11/4}\sqrt{N_{p}}}\left(G\mathcal{M}\right)^{5/4}\frac{L}{D_{L}}\times\left(\sqrt{t_{c}}-\sqrt{t_{c}-T_{\rm obs}}\right)^{1/2}$ (17)
$\approx$ $8.4\left(\frac{\mathcal{M}}{100~{}M_{\odot}}\right)^{5/4}\frac{100~{}{\rm Mpc}}{D_{L}}\times\left(\sqrt{\frac{t_{c}}{\rm 1~{}month}}-\sqrt{\frac{t_{c}-T_{\rm obs}}{\rm 1~{}month}}\right)^{1/2},$
@@
$C$ $=$ $-\frac{0.388101}{c^{13/3}N_{p}}\left(G\mathcal{M}\right)^{5/3}\frac{L^{10/3}}{D_{L}^{2}}\left[\mathcal{F}(t_{c})-\mathcal{F}(t_{c}-T_{\rm obs})\right],$ (18)
$\approx$ $162\left(\frac{\mathcal{M}}{100~{}M_{\odot}}\right)^{5/3}\left(\frac{100~{}{\rm Mpc}}{D_{L}}\right)^{2}\left[\mathcal{F}(t_{c})-\mathcal{F}(t_{c}-T_{\rm obs})\right],$
@@
$\mathcal{F}(t)$ $=$ ${\rm arctan}\left[0.57735-\frac{0.577948~{}c^{7/12}L^{2/3}}{\left(G\mathcal{M}\right)^{5/12}t^{1/4}}\right],$ (19)
$\approx$ ${\rm arctan}\left[0.57735-\frac{0.23866}{\left(\frac{\mathcal{M}}{100~{}M_{\odot}}\right)^{5/12}\left(\frac{t}{\rm 1~{}month}\right)^{1/4}}\right]$
In our test, as shown in Figure 5, we found that in the case of ASA, our formula provided good estimation accuracy for SBBHs. Without incorporating the lowest-order correction, we could make good estimates of the SNR for signals ranging from $7\sim 400~{}M_{\odot}$. The SNR was overestimated for signals outside this range. When the lowest-order correction was applied, the SNR of most signals with masses below $400~{}M_{\odot}$ could be accurately calculated.
@@
Despite our formula providing good estimations in the case of ASA, the SBBH signal with response is challenging to compute analytically. As illustrated in Figure 6, after considering the TDI response, the low-frequency component of the PSD remains smooth, allowing the SNR of the MBHB signal with response to be analytically approximated using LFA. While in the high-frequency regime where SBBH resides, the PSD becomes highly complex and lacks a satisfactory formula for approximation. However, it is important to note that while we consider a three-month observation for all systems, SBBHs can keep inspiraling within the sensitive bands of space detectors for years. Therefore, our estimation of the ASA SNR will also be applicable for such long-duration inspiral-only observations.
@@
5 Conclusion
In this work, we present a simplified analytic formula for the SNR of BBH systems in TianQin. The analytical formula for low-frequency scenarios is applicable to signals from MBHBs with chirp masses above approximately $2\times 10^{3}~{}M_{\odot}$, while the analytical formula for high-frequency scenarios is suitable for signals from SBBHs with chirp masses below approximately $400~{}M_{\odot}$. In the Table 1, we summarize all the formulas, their applicability, and the assumptions in this paper.
Table 1: The table summarize all the formulas, their applicability, and the assumptions in this paper.
Sources SNR formulae Applicability($M_{\odot}$)1 Assumptions
MBHB Zero-order formulae (7) $2\times 10^{3}\sim 5\times 10^{4}$ LFA(31), $f$-independent $S_{a}$(31)
Lowest-order formulae (10,11) $2\times 10^{3}\sim 2\times 10^{5}$ LFA(31), Integral expansion(34)
Response factor (13,15) $\geq 2\times 10^{3}$ LFA(37,40,46), $P_{\rm res}$ factorization 15
SBBH Zero-order formulae (17) $7\sim 400$ SWA(16)
Lowest-order formulae (18) $1\sim 400$ SWA(16)
1The applicability indicates the mass range where formula error stays below 10% one month prior to merger. Except near merger, the applicability shows little variation over time.
On the one hand, our findings can enable researchers in various fields to quickly and accurately estimate the SNR of GW signals from different types of BBH systems. On the other hand, our results reveal the accumulation pattern of the SNR for BBH signals, which can be applied to the identification of future signals and the construction of detection statistics.
We evaluated the significance of errors resulting from using SNR analytical formulae. For TianQin, our formula of MBHB can accurately estimate the inspiral signals within the mass range of $2\times 10^{3}\sim 5\times 10^{4}~{}M_{\odot}$. Incorporating a lowest-order correction can extend the range of accurate estimates between $2\times 10^{3}\sim 2\times 10^{5}~{}M_{\odot}$. However, for LISA, achieving good estimation accuracy is challenging, regardless of whether the lowest-order correction is considered.
We derived an accurate analytical formula for the SNR of SBBHs in the ASA case. The formula can accurately estimate the inspiral signals within the mass range of $7$ to $400~{}M_{\odot}$. Upon the application of the lowest-order correction, the SNR of the majority of signals with chirp mass below $400~{}M_{\odot}$ could be accurately calculated. Due to the complexity of the PSD at high frequencies, it remains challenging to provide a precise analytical formula for the SNR of SBBH signals with response. We will delve into this scenario in the future.